# A Systematic Digital Engineering Approach to Verification & Validation of Autonomous Ground Vehicles in Off-Road Environments

### v1

@@ 1 Introduction
  • Integrating the digital twin framework with MBSE and MBD workflows to recursively design and test off-road autonomy algorithms by developing an open interface for MathWorks System Composer [^2] https://www.mathworks.com/products/system-composer.html.
@@ 3 Digital Engineering Framework
This section highlights the proposed digital engineering framework for conceptualizing, designing, developing, integrating, and validating off-road autonomy algorithms. The framework couples the high-fidelity digital twin simulations of AutoDRIVE Simulator Samak et al. (2021) with an integrated model-based engineering (MBSE + MBD) approach in MathWorks System Composer, which allows systematic development and assessment of off-road autonomy algorithms.
@@ 3.1.1 Vehicle Simulation
The vehicle’s powertrain generates a torque equivalent to $\tau_{\text{total}}=\left[\tau_{e}\right]_{RPM_{e}}*\left[GR\right]_{G_{\#}}*FDR*\tau*\mathscr{A}$, were, $\tau_{e}$ is the engine torque, $\tau$ is the throttle input, and $\mathscr{A}$ is a non-linear smoothing operator. The engine speed smoothly updates as $RPM_{e}:=\left[RPM_{i}+\left(|RPM_{w}|*FDR*GR\right)\right]_{(RPM_{e},v)}$ where, $RPM_{i}$ is the idle speed, $RPM_{w}$ is the mean wheel speed, $FDR$ is the final drive ratio, $GR$ is the current gear ratio, and $v$ is the vehicle velocity. It is kept in check by the automatic transmission, which maintains a good operating range for a given speed: $RPM_{e}=\frac{{v_{\text{MPH}}*5280*12}}{{60*2*\pi*R_{\text{tire}}}}*FDR*GR$. The output torque at wheels depends on the vehicle’s drive configuration: $\tau_{\text{out}}=\begin{cases}\frac{\tau_{\text{total}}}{2}&\text{if FWD/RWD}\\ \frac{\tau_{\text{total}}}{4}&\text{if AWD}\end{cases}$. The actual torque transmitted to the wheels $\tau_{w}$ is modeled by dividing the output torque $\tau_{\text{out}}$ to the left ${}^{L}\tau_{w}=\tau_{\text{out}}*(1-\tau_{\text{drop}}*|\delta^{-}|)$ and right ${}^{R}\tau_{w}=\tau_{\text{out}}*(1-\tau_{\text{drop}}*|\delta^{+}|)$ wheels based on the steering input. Here, $\tau_{\text{drop}}$ is the torque drop at differential and $(\tau_{\text{drop}}*|\delta^{\pm}|)$ is clamped between $[0,0.9]$.
@@
The wheel-tire interconnect is simulated by modeling the friction curve for each tire, represented as $\left\{\begin{matrix}{{}^{i}F_{t_{x}}}=F(^{i}S_{x})\\ {{}^{i}F_{t_{y}}}=F(^{i}S_{y})\\ \end{matrix}\right.$, where ${}^{i}S_{x}$ and ${}^{i}S_{y}$ denote the longitudinal and lateral slip values of the $i$-th tire. Here, the friction curve is modeled by a two-piece cubic spline $F(S)=\left\{\begin{matrix}f_{0}(S);\;\;S_{0}\leq S<S_{e}\\ f_{1}(S);\;\;S_{e}\leq S<S_{a}\\ \end{matrix}\right.$, with $f_{k}(S)=a_{k}*S^{3}+b_{k}*S^{2}+c_{k}*S+d_{k}$. The first segment of the spline spans from the origin $(S_{0},F_{0})$ to the extremum $(S_{e},F_{e})$, while the second segment ranges from the extremum $(S_{e},F_{e})$ to the asymptote $(S_{a},F_{a})$.
@@ 3.1.2 Sensor Simulation
The virtual cameras are simulated using a standard viewport rendering pipeline. First, the camera view matrix $\mathbf{V}\in SE(3)$ is computed by obtaining the relative homogeneous transform of the camera $\{c\}$ with respect to the world $\{w\}$. Next, the camera projection matrix $\mathbf{P}\in\mathbb{R}^{4\times 4}$ is calculated to project world coordinates into image space coordinates: $\mathbf{P}=\begin{bmatrix}\frac{2*N}{R-L}&0&\frac{R+L}{R-L}&0\\ 0&\frac{2*N}{T-B}&\frac{T+B}{T-B}&0\\ 0&0&-\frac{F+N}{F-N}&-\frac{2*F*N}{F-N}\\ 0&0&-1&0\\ \end{bmatrix}$, where $N$ and $F$ are the near and far clipping planes of the camera, while $L$, $R$, $T$, and $B$ denote the left, right, top, and bottom offsets of the sensor. Finally, a post-processing step replicates physical phenomena like lens distortion, depth of field, exposure, ambient occlusion, contact shadows, bloom, motion blur, film grain, chromatic aberration, and other similar lens and film effects.
@@ 3.1.4 Application Programming Interface
In order to facilitate bi-directional data flow between the digital twins and the system under test (SUT), an application programming interface (API) is necessary. Consequently, a significant effort went into developing an object-oriented WebSocket protocol to establish a direct, real-time communication bridge between AutoDRIVE and MathWorks tools. The base API is directly compatible with MATLAB scripts, and can be wrapped as Level-2 MATLAB S-Functions to create custom blocks that seamlessly integrate the API with Simulink models. The said API can be abstracted and exposed to enable granular control of the vehicles (sensor data streaming, actuator/light control, etc.) as well as their operating environments (time of day, weather, etc.).
@@ 4.0.3 Requirements Engineering
Table 1 hosts the requirements imposed on the perception (R1), planning (R2), and control (R3) sub-systems as well as those (R4) imposed on the system as a whole.
@@ 4.0.6 Test Results
We further demonstrate how verification scores across different variants can quickly inform design choices with credibility. For the perception sub-system (C1), we can observe that the R1 satisfaction rate of C1.2 is significantly higher than that of C1.1. Furthermore, its overall satisfaction rate is also higher, making it the better design choice. Similarly, for the planning sub-system, although the R2 satisfaction rate of C2.1 and C2.2 is almost the same, the overall satisfaction with C2.1 is over two times higher, giving it a competitive edge in terms of reliability. Finally, for the control sub-system, we can see that C3.2 dominates across all requirements, except for R4. Consequently, it can be established that while C3.2 is better in terms of its tracking and comfort, C3.1 is much better in terms of safety.
Additionally, the verification scores across different parameters can help reliably define the operational design domain of the SUT. We can observe that the SUT did not perform well in the evening/night, which can be primarily attributed to the perception sub-system. None of the test cases during these times passed all the requirements. Consequently, the SUT may target only morning and afternoon times of the day, which highlight perfect scores for R1. A similar comparison among various weather conditions reveals that the SUT performs equally across all of them, with slightly less reliability during snowfall. Consequently, the SUT may target all weather conditions, with extra caution when it is snowing.

### v2

@@ 1 Introduction
  • Integrating the digital twin framework with MBSE and MBD workflows to recursively design and test off-road autonomy algorithms by developing an open interface for MathWorks® System Composer [^2] https://www.mathworks.com/products/system-composer.html.
@@ 3 Digital Engineering Framework
This section highlights the proposed digital engineering framework for conceptualizing, designing, developing, integrating, and validating off-road autonomy algorithms. The framework couples the high-fidelity digital twin simulations of AutoDRIVE Simulator Samak et al. (2021) with an integrated model-based engineering (MBSE + MBD) approach in MathWorks® System Composer, which allows systematic development and assessment of off-road autonomy algorithms.
@@ 3.1.1 Vehicle Simulation
The vehicle’s powertrain generates a torque equivalent to $\tau_{\text{total}}=\left[\tau_{e}\right]_{RPM_{e}}*\left[GR\right]_{G_{\#}}*FDR*\tau*\mathscr{A}$, were, $\tau_{e}$ is the engine torque, $\tau$ is the throttle input, and $\mathscr{A}$ is a non-linear smoothing operator. The engine speed smoothly updates as $RPM_{e}:=\left[RPM_{i}+\left(|RPM_{w}|*FDR*GR\right)\right]_{(RPM_{e},v)}$ where, $RPM_{i}$ is the idle speed, $RPM_{w}$ is the mean wheel speed, $FDR$ is the final drive ratio, $GR$ is the current gear ratio, and $v$ is the vehicle velocity. It is kept in check by the automatic transmission, which maintains a good operating range for a given speed: $RPM_{e}=\frac{{v_{\text{MPH}}*5280*12}}{{60*2*\pi*R_{\text{tire}}}}*FDR*GR$. The output torque at wheels depends on the vehicle’s drive configuration: $\tau_{\text{out}}=\begin{cases}\frac{\tau_{\text{total}}}{2}&\text{if FWD/RWD}\\ \frac{\tau_{\text{total}}}{4}&\text{if AWD}\end{cases}$. The actual torque transmitted to the wheels $\tau_{w}$ is modeled by dividing the output torque $\tau_{\text{out}}$ to the left ${}^{L}\tau_{w}=\tau_{\text{out}}*(1-\tau_{\text{drop}}*|\delta^{-}|)$ and right ${}^{R}\tau_{w}=\tau_{\text{out}}*(1-\tau_{\text{drop}}*|\delta^{+}|)$ wheels based on the steering input. Here, $\tau_{\text{drop}}$ is the torque drop at differential and $(\tau_{\text{drop}}*|\delta^{\pm}|)$ is clamped between [0,0.9][0,0.9][ 0 , 0.9 ].
@@
The road-tire interconnect is simulated by modeling the friction curve for each tire, represented as $\left\{\begin{matrix}{{}^{i}F_{t_{x}}}=F(^{i}S_{x})\\ {{}^{i}F_{t_{y}}}=F(^{i}S_{y})\\ \end{matrix}\right.$, where ${}^{i}S_{x}$ and ${}^{i}S_{y}$ denote the longitudinal and lateral slip values of the $i$-th tire. Here, the friction curve is modeled by a two-piece cubic spline $F(S)=\left\{\begin{matrix}f_{0}(S);\;\;S_{0}\leq S<S_{e}\\ f_{1}(S);\;\;S_{e}\leq S<S_{a}\\ \end{matrix}\right.$, with $f_{k}(S)=a_{k}*S^{3}+b_{k}*S^{2}+c_{k}*S+d_{k}$. The first segment of the spline spans from the origin $(S_{0},F_{0})$ to the extremum $(S_{e},F_{e})$, while the second segment ranges from the extremum $(S_{e},F_{e})$ to the asymptote $(S_{a},F_{a})$.
@@ 3.1.2 Sensor Simulation
The virtual cameras are simulated using a standard viewport rendering pipeline. First, the camera view matrix $\mathbf{V}\in SE(3)$ is computed by obtaining the relative homogeneous transform of the camera $\{c\}$ w.r.t. world $\{w\}$. Next, the camera projection matrix $\mathbf{P}\in\mathbb{R}^{4\times 4}$ is calculated to project world coordinates into image space coordinates: $\mathbf{P}=\begin{bmatrix}\frac{2*N}{R-L}&0&\frac{R+L}{R-L}&0\\ 0&\frac{2*N}{T-B}&\frac{T+B}{T-B}&0\\ 0&0&-\frac{F+N}{F-N}&-\frac{2*F*N}{F-N}\\ 0&0&-1&0\\ \end{bmatrix}$, where $N$ and $F$ are the near and far clipping planes of the camera, while $L$, $R$, $T$, and $B$ denote the left, right, top, and bottom offsets of the sensor. Finally, a post-processing step replicates physical phenomena like lens distortion, depth of field, exposure, ambient occlusion, contact shadows, bloom, motion blur, film grain, chromatic aberration, and other lens and film effects.
@@ 3.1.4 Application Programming Interface
In order to facilitate bi-directional data flow between the digital twins and the system under test (SUT), an application programming interface (API) is necessary. Consequently, a significant effort went into developing an object-oriented WebSocket protocol to establish a direct, real-time communication bridge between AutoDRIVE and MathWorks® tools. The base API is directly compatible with MATLAB® scripts, and can be wrapped as Level-2 MATLAB® S-Functions to create custom blocks that seamlessly integrate the API with Simulink models. The said API can be abstracted and exposed to enable granular control of the vehicles (sensor data streaming, actuator/light control, etc.) and their operating environments (time of day, weather, etc.).
@@ 4.0.3 Requirements Engineering
Table 1 hosts the requirements imposed on the perception (R1), planning (R2), and control (R3) sub-systems, as well as those imposed on the system (R4) as a whole.
@@ 4.0.6 Test Results
We further demonstrate how verification scores across different variants can quickly inform design choices with credibility. For the perception sub-system, we can observe that the R1 satisfaction rate of C1.2 is significantly higher than that of C1.1. Furthermore, its overall satisfaction rate is also higher, making it a better design choice. Similarly, for the planning sub-system, although the R2 satisfaction rate of C2.1 and C2.2 is almost the same, the overall satisfaction with C2.1 is over two times higher, giving it a competitive edge in terms of reliability. Finally, for the control sub-system, we can see that C3.2 dominates across all requirements, except for R4. Consequently, it can be established that while C3.2 is better in terms of tracking and comfort, C3.1 is much better in terms of safety.
Finally, the verification scores across different parameters can help reliably define the operational design domain of the SUT. We can observe that the SUT did not perform well in the evening/night, which can be primarily attributed to the perception sub-system. None of the test cases during these times passed all the requirements. Consequently, the SUT may target only morning and afternoon times of the day, which highlight perfect scores for R1. A similar comparison among various weather conditions reveals that the SUT performs equally across all of them, with slightly less reliability during snowfall. Consequently, the SUT may target all weather conditions, with extra caution when it is snowing.